OpenStudy (anonymous):

consider a swuare with side of lengh 's' and diagonal of length 'd'. write a formula expressing 'd' as a function of 's'...............is this correct: d(s) ?

OpenStudy (anonymous):

\[d=s\sqrt{2}\] This is because: \[d^2=s^2+s^2\] \[d^2=2s^2\] Take square of both sides: \[d=\sqrt{2s^2}\] Simplify: \[d(s)=s\sqrt{2}\]